# DeLTa: Demonstration and Language-Guided Novel Transparent Object Manipulation

### v1

@@ Abstract
Despite the prevalence of transparent object interactions in human everyday life, transparent robotic manipulation research remains limited to short-horizon tasks and basic grasping capabilities. Although some methods have partially addressed these issues, most of them have limitations in generalizability to novel objects and are insufficient for precise long-horizon robot manipulation. To address this limitation, we propose DeLTa (Demonstration and Language-Guided Novel Transparent Object Manipulation), a novel framework that integrates depth estimation, 6D pose estimation, and vision-language planning for precise long-horizon manipulation of transparent objects guided by natural task instructions. A key advantage of our method is its single-demonstration approach, which generalizes 6D trajectories to novel transparent objects without requiring category-level priors or additional training. Additionally, we present a task planner that refines the VLM-generated plan to account for the constraints of a single-arm, eye-in-hand robot for long-horizon object manipulation tasks. Through comprehensive evaluation, we demonstrate that our method significantly outperforms existing transparent object manipulation approaches, particularly in long-horizon scenarios requiring precise manipulation capabilities. Project page: https://sites.google.com/view/DeLTa25/
@@ I INTRODUCTION
Transparent object pose estimation methods [9, 10, 11, 12] often adopt category prior knowledge to estimate poses of novel object instances within the same category. As a result, robotic manipulation for transparent objects is inherently restricted to category-level object pose estimation [13, 14, 15]. While category-level pose approaches have achieved promising results in generalizing to unknown objects within the same category, they struggle to generalize to novel objects beyond their trained categories. Moreover, their disregard of fine-grained object geometry limits applications to precise manipulation under certain task constraints (e.g., align beverage bottles in a straight row when stocking a grocery shelf) even for novel object instances within the same trained category. This makes novel object instance pose estimation methods [5, 16, 17] more desirable in such scenarios.
@@
  • $\bullet$
    We show that a single human demonstration per primitive skill suffices to generalize to novel objects, with the demonstration trajectory guiding obstacle-aware robot motion planning.
  • $\bullet$
    We propose a VLM-guided planner that decomposes natural language instructions into multi-step actions, refines them with validation and search to enforce robot-specific constraints (e.g., one-armed, eye-in-hand), and retrieves object meshes and demonstration trajectories from pre-computed databases for transparent object manipulation.
@@ II Related Works
II-A Transparent Object Manipulation
Most transparent object manipulation research has primarily focused on short-horizon grasping tasks, using either multiple views [30, 31, 32, 33], single view [34, 35, 36, 37]. Most existing methods focus on reconstructing depth for grasping and have limited capabilities for long-horizon tasks that require precise manipulation from instructions.
II-B Transparent Object Pose Estimation
One straightforward approach for achieving precise manipulation is through 6D object pose estimation, which suits robots operating in 3D space [5, 38, 39]. Most object pose estimation research [5, 16] has focused on non-transparent objects. While recent works [11, 10, 9] have addressed transparent object pose estimation, these methods remain limited to category-level understanding and still struggle to generalize to novel instances. Novel object pose estimation for transparent objects remains a challenging and open problem.
II-C From Human Demonstration to Robot Skills
Teaching robots to perform human tasks necessitates intuitive and efficient ways that operate without requiring wearable sensors [25, 26] or teleoperation [24]. The most natural approach involves demonstrating a task once and enabling the robot to replicate the observed behavior [20, 40]. However, these methods face significant limitations in transparent object perception and long-horizon manipulation scenarios, as they rely on traditional RGB-D perception and primarily focus on short-term tasks. Moreover, these demonstration-based methods generally lack obstacle-avoidance capability [41], a critical requirement for real-world deployment.
II-D VLM-guided Long-Horizon Task Planning
While significant progress has been made in applying VLMs to robotic long-horizon task planning [42, 43], current approaches are often limited to simplified tasks or simulation environments and often assume complete prior knowledge of environmental objects [44, 45], including ground-truth object poses. For instance, [43] repeatedly queries LLMs to generate symbolic plans but does not handle perception errors or real-time execution. This leaves a critical gap between noisy perception, motion planning, and symbolic planning, which occurs even more frequently for transparent object manipulation. Therefore, a unified framework for transparent object manipulation addressing these limitations is highly desirable.
@@ III Method
Our goal is to enable robots to execute manipulation tasks on novel transparent objects by leveraging single-object human demonstration trajectories from task instructions.
@@
We first explain our approach for parsing single-object demonstration trajectories from human videos (Sec. III-A). We then describe the task planning process for long-horizon tasks from natural language instructions (Sec. III-B). Finally, we detail how these trajectories are transferred to novel objects during robotic manipulation (Sec. III-C).
@@ III-A Parsing Human Video Demonstration
Fig. 2 (top-left) illustrates our human demonstration parsing pipeline, which extracts single-object demonstration trajectories from video demonstrations. In this work, we mainly consider three skill primitives: pick, place, and pour. For each skill primitive, we extract its trajectory from a single video demonstration of a randomly selected object. These extracted trajectories are then stored as trajectory database. During robot execution, the action manager selects the trajectory-based skill corresponding to the primitive action, a basic task unit used by the task planner to compose high-level plans (Sec. III-C). Notably, our method requires only a single demonstration per skill, enabling cross-object transfer to manipulate novel objects. This is in contrast to multiple separate trajectories for each target object, typically required as in prior works [21, 40].
@@
Transparent Depth Estimation. Fig. 3 shows the challenge that raw sensor depth from ZED stereo camera [2] fails to capture the surface depth of transparent objects. To address this limitation, we harness FoundationStereo [46], a foundation model for stereo depth estimation. It takes stereo images from the ZED camera as input and outputs pixel-wise metric-scale depth. The reconstructed depth enhances the overall robot manipulation pipeline by improving object and hand pose estimation, as well as high-quality 3D collision map reconstruction for the motion planner.
@@
Mesh Database. We pre-build object mesh databases containing textured object shapes to be used for object pose estimation during both demonstration video parsing and robot execution. Meshes are obtained via image-to-3D reconstruction [50, 51] and existing databases [52], which are then transformed to the coordinate frame of the estimated hand pose (as will be discussed in the following). This database contains geometry information with textures, serving exclusively for robot execution pose estimation.
@@ III-B Vision-Language Guided Task Planning
The task planner takes a high-level task instruction as input and generates a task plan consisting of primitive actions (e.g., Pick, Place, PlaceInFront, LookFor) for the robot to execute. It consists of three main components: VLM planner, plan translator, and plan grounding (Fig. 2 bottom-left). The VLM planner processes task instructions and visual inputs to produce an initial high-level plan. The plan translator converts this plan into the formal planning language PDDL [54], [55] while checking for syntactic errors. Finally, the plan grounding takes into account robot constraints (e.g., limited FoV under eye-in-hand camera configuration), evaluates feasibility, and adds intermediate actions (e.g., look for objects outside view) to ensure executability through an iterative search-and-refinement process.
@@
VLM Planner. The VLM outputs a sequence of actions $\pi_{VLM}$, where each action consists of an action type (e.g., Pick) and a target object or location. As shown in Fig. 4(a), the prompt specifies the task goal, known locations, objects, and available actions, along with representation constraints and a fixed format to ensure a clear and ordered action plan. This sequence serves as a skeleton plan [44], which is subsequently grounded and refined. We utilize a foundation VLM [56] without fine-tuning to preserve generalization and maintain robust reasoning across diverse tasks.
@@
We divide $\pi_{VLM}$ into subtasks $\Pi=\{\pi^{1},\dots,\pi^{N}\}$ at placement actions, ensuring free robot hands for independent subtask execution. Each subtask labels object manipulation actions (e.g., Pick) as key actions $A_{k}$ and others (e.g., LookFor) as connecting actions $A_{c}$. Key actions are validated sequentially; if an action fails, a backward breadth-first search is triggered to find satisfying predecessor sequences using $A_{c}$ and the order-independent primitive actions $A_{s}$. This process includes actions to move the robot joints into suitable configurations when they are not appropriate for executing the next action. While performing the search, it tracks the visited plans $V$ and records used actions $U$. Search failure returns unsatisfied preconditions and partial plans to the VLM for replanning.
@@
Plan Refinement bridges VLM reasoning and robot execution through iterative feedback. Refinement occurs in two cases: (1) PDDL syntax errors during translation, generating format correction messages like ”Failed to create {action} instance: {error}.”; and (2) task planning errors during search when connecting actions cannot be found, prompting precondition error messages with partial plans. This process iterates until producing a fully grounded, executable plan, as shown in Fig. 2.
@@ III-C Demonstration-Guided Robot Action Execution
The robot action execution stage implements the action sequence generated in Sec. III-B through the trajectory and mesh database from Sec. III-A. This execution framework integrates three core components: (1) target object pose estimation, (2) target trajectory generation, and (3) collision-aware motion planning.
Action Manager sequentially processes planner-generated actions by selecting target meshes and primitive skill, generating motions, and evaluating feasibility and termination. Target mesh selection maps VLM action parameters (e.g., cola_on_the_shelf) to corresponding mesh database objects (e.g., cola) using an LLM with simple prompts, as shown in Fig. 4(b). Action feasibility is evaluated based on predefined precondition that depend on the current robot and environment states. Upon action completion, the action manager notifies the task planner to dispatch the next action.
@@
During the second stage, the robot tracks $^{B}\tau_{\mathrm{al}}$ using the same QP-based IK solver with adaptive accuracy. We set different tolerances for the solver depending on trajectory progress. For the first $\alpha$ of the trajectory, a larger tolerance is applied to provide flexibility for collision avoidance and joint limit handling. In the final $1-\alpha$, a smaller tolerance is enforced to ensure precise manipulation near the target object. This hierarchical strategy achieves both safe obstacle avoidance and high-precision execution during the final phase of the task. The framework is insensitive to the choice of $\alpha$, which is set to 0.8 empirically.
@@ IV Experiments
The evaluation is conducted through real-world experiments using a Kinova Gen3 7-DOF arm equipped with a Robotis RH-P12-RN gripper and an eye-in-hand ZED stereo camera. For evaluation, 9 different transparent and non-transparent objects were used, as illustrated in Fig. 6. Fig. 7 illustrates the sequence of three manipulation tasks considered for evaluation:
@@
Task2: Chemical Experiment. ”Can you make [target color] liquid in the [target object]?” This task simulates laboratory scenarios using transparent objects filled with various liquids. We tested with seven transparent objects and two target colors, requiring long-horizon planning, precise pouring, and collision avoidance in dense environments.
Task3: Grocery Stocking. ”Arrange the [objects] on the staging area to the shelf in a straight row as shown in the reference image, placing each one directly in front of the previous one.” Inspired by retail scenarios, this automatic organization task requires long-horizon planning, collision avoidance, and precise pose estimation for aligned stocking.
For each task, we conducted 10 trials with different object instances and pose variations. The three single-object trajectories (pick, place, pour) were extracted from the object shown in the top left of Fig. 6. Our framework supports 10 primitive actions (LookForAt, LookFor, Pick, Pour, PlaceBack, Place, PlaceBetween, PlaceInFront, Face, InitPose). The environments include a shelf (Task 1), a laboratory area (Task 2), and a staging area with a shelf (Task 3).
@@ IV-A Baselines
 Method Success Rate (%) ↑
Task 1 Task 2 Task 3
ClearGrasp [18] 70 0 20
YODO [20] 70 0 0
[…]
Table I: Comparison with state-of-the-art methods.
@@ IV-B Real World Results
For each task, we evaluated performance using task success rate, as reported in table I. We observed that all methods perform well on Tight Shelf Retrieval (Task 1), since most existing methods focus primarily on this type of task. However, for long-horizon tasks involving precise manipulation—such as chemical experiment and grocery stocking—the performance of YODO [20] and ClearGrasp [18] drops significantly.
@@
Depth 6D Pose
Plan
Refinement
Plan
Search
Task 1 Task 2 Task 3
@@
Table II: Ablation study of our core components on three different tasks in real world.
IV-C Ablation Studies
(1) Is depth estimation necessary for transparent object manipulation? To answer this question, we evaluate our method using raw depth sensor observations instead of our depth estimation results (Fig. 3). The comparison between Table II-(1) and Table II-Ours demonstrates that our depth estimation significantly improves performance across all tasks. Specifically, in the grocery stocking task, inaccurate depth information hindered the robot from compactly arranging objects within the shelf dimensions. This verifies the significant benefits of enhanced depth estimation for precise transparent robotics manipulation, as raw depth sensors alone are insufficient for handling transparent objects.
(2) Is 6D pose estimation necessary compared to 3D position-based manipulation? One of the straightforward approaches for robot manipulation uses 3D position from open vocabulary segmentation and projects to 3D using depth for manipulation. As shown in Table II-(2), this position-based method achieves reasonable performance in Tight Shelf Retrieval (70%) and Chemical Experiment (60%) tasks, but fails in grocery stocking tasks, demonstrating the limitations of position-only approaches for grasping diverse rotation. Specifically, when objects are lying with small rotation errors, grasps often fail, highlighting the necessity of accurate 6D pose for robust manipulation.
(3-4) Can VLMs effectively handle long-horizon task planning? While naive VLMs may appear capable of long-horizon task planning, LLMs lack the ability to incorporate robot manipulation constraints into action sequences For example, in a pick-and-place task, the model may attempt to place an object before actually grasping it. Table II-(3) shows that a naive LLM mostly fails to generate valid plans, even for a short-horizon task such as tight shelf retrieval. An alternative approach is to iteratively refine the plan using a validator and repeated queries with error feedback, as in prior studies [42]. However, Table II-(4) demonstrates that this approach still struggles with long-horizon tasks, such as chemical experiments and grocery stocking, when the number of refinement iterations is limited to 10. Even when task planning succeeds, execution failures can still occur during the motion planning stage because this method relies solely on symbolic states without incorporating continuous variables such as joint poses, unlike our search-based method. Compared to Table II-(3-4) of VLM planning, our proposed search-based refinement successfully handles long-horizon planning.

### v2

@@ Abstract
Despite the prevalence of transparent object interactions in human everyday life, transparent robotic manipulation research remains limited to short-horizon tasks and basic grasping capabilities. Although some methods have partially addressed these issues, most of them have limitations in generalization to novel objects and are insufficient for precise long-horizon robot manipulation. To address this limitation, we propose DeLTa (Demonstration and Language-Guided Novel Transparent Object Manipulation), a novel framework that integrates depth estimation, 6D pose estimation, and vision-language planning for precise long-horizon manipulation of transparent objects guided by natural language task instructions. A key advantage of our method is its single-demonstration approach, which generalizes 6D trajectories to novel transparent objects without requiring category-level priors or additional training. Additionally, we present a task planner that refines the VLM-generated plan to account for the constraints of a single-arm, eye-in-hand robot for long-horizon object manipulation tasks. Through comprehensive evaluation, we demonstrate that our method significantly outperforms existing transparent object manipulation approaches, particularly in long-horizon scenarios requiring precise manipulation capabilities. Project page: https://sites.google.com/view/DeLTa25/
@@ I INTRODUCTION
Transparent object pose estimation methods [37, 58, 4, 30] often adopt category prior knowledge to estimate poses of novel object instances within the same category. As a result, robotic manipulation for transparent objects is inherently restricted to category-level object pose estimation [25, 16, 24]. While category-level pose approaches have achieved promising results in generalizing to unknown objects within the same category, they struggle to generalize to novel objects beyond their trained categories. Moreover, their disregard of fine-grained object geometry limits applications to precise manipulation under certain task constraints (e.g. align beverage bottles in a straight row when stocking a grocery shelf) even for novel object instances within the same trained category. This makes novel object instance pose estimation methods [53, 23, 26] more desirable in such scenarios.
@@
  • We show that by combining pose-based trajectory retargeting with canonicalized object meshes, a single human demonstration per primitive skill can be transferred to novel objects, eliminating the need for per-object demonstrations.
  • We propose a VLM-guided planner that decomposes natural language instructions into multi-step actions, refines them with validation and search to enforce robot-specific constraints (e.g., single-arm, eye-in-hand), and retrieves object meshes and demonstration trajectories from pre-computed databases for transparent object manipulation.
@@ II Related Works
Transparent Object Manipulation: Most transparent object manipulation research has primarily focused on short-horizon grasping tasks, using either multiple views [15, 7, 22, 18], single view [9, 45, 17, 49]. Most existing methods focus on reconstructing depth for grasping and have limited capabilities for long-horizon tasks that require precise manipulation from instructions.
Transparent Object Pose Estimation: One straightforward approach for achieving precise manipulation is through 6D object pose estimation, which suits robots operating in 3D space [53, 47, 28]. Most object pose estimation research [53, 23] has focused on non-transparent objects. While recent works [4, 58, 37] have addressed transparent object pose estimation, these methods remain limited to category-level understanding and still struggle to generalize to novel instances. Novel object pose estimation for transparent objects remains a challenging and open problem.
From Human Demonstration to Robot Skills: Teaching robots to perform human tasks necessitates intuitive and efficient ways that operate without requiring wearable sensors [48, 5] or teleoperation [46]. The most natural approach involves demonstrating a task once and enabling the robot to replicate the observed behavior [52, 32]. However, these methods face significant limitations in transparent object perception and long-horizon manipulation scenarios, as they rely on traditional RGB-D perception and primarily focus on short-term tasks. Moreover, these demonstration-based methods generally lack obstacle-avoidance capability [50], a critical requirement for real-world deployment.
VLM-guided Long-Horizon Task Planning: While significant progress has been made in applying VLMs to robotic long-horizon task planning [60, 6], current approaches are often limited to simplified tasks or simulation environments and often assume privileged information of environmental objects [56, 59], including ground-truth object poses. For instance, [6] repeatedly queries LLMs to generate symbolic plans but does not handle perception errors or real-time execution. This leaves a critical gap between noisy perception, motion planning, and symbolic planning, which occurs even more frequently for transparent object manipulation. Therefore, a unified framework for transparent object manipulation addressing these limitations is highly desirable.
@@ III Method
Our goal is to enable robots to execute manipulation tasks on novel transparent objects, given a task instruction, by leveraging single-object human demonstration trajectories.
@@
We first explain our approach for parsing single-object demonstration trajectories from human videos (Section III-A). We then describe the task planning process for long-horizon tasks from natural language instructions (Section III-B). Finally, we detail how these trajectories are transferred to novel objects during robotic manipulation (Section III-C).
@@ III-A Parsing Human Video Demonstration
Fig. 2 (top-left) illustrates our human demonstration parsing pipeline, which extracts single-object demonstration trajectories from video demonstrations. In this work, we mainly consider three skill primitives: pick, place, and pour. For each skill primitive, we extract its trajectory from a single video demonstration of a randomly selected object. These extracted trajectories are then stored as trajectory database. During robot execution, the action manager selects the trajectory-based skill corresponding to the primitive action, a basic task unit used by the task planner to compose high-level plans (Sec. III-C). Notably, our method requires only a single demonstration per skill, enabling cross-object transfer to manipulate novel objects. This is in contrast to multiple separate trajectories for each target object, required as in prior works [13, 32].
@@
Transparent Depth Estimation. Fig. 3 shows the challenge that raw sensor depth from ZED stereo camera [43] fails to capture the surface depth of transparent objects. To address this limitation, we harness FoundationStereo [54, 51], a foundation model for stereo depth estimation. It takes stereo images from the ZED camera as input and outputs pixel-wise metric-scale depth. The reconstructed depth enhances the overall robot manipulation pipeline by improving object and hand pose estimation, as well as high-quality 3D collision map reconstruction for the motion planner.
@@
Mesh Database. We pre-build object mesh databases containing textured object shapes to be used for object pose estimation during both demonstration video parsing and robot execution. Meshes are obtained via image-to-3D reconstruction [55, 14] and existing databases [21]. Each mesh is canonicalized and aligned to the coordinate frame of the estimated hand pose from the demonstration video, ensuring consistent trajectory retargeting across novel objects. This database contains geometry information with textures, serving exclusively for robot execution pose estimation.
@@ III-B Vision-Language Guided Task Planning
The task planner takes a high-level task instruction as input and generates a task plan consisting of primitive actions (e.g. Pick, Place, PlaceInFront, LookFor) for the robot to execute. It consists of three main components: VLM planner, plan translator, and plan grounding (Fig. 2 bottom-left). The VLM planner processes task instructions and visual inputs to produce an initial high-level plan. The plan translator converts this plan into the formal planning language PDDL [10], [12] while checking for syntactic errors. Finally, the plan grounding takes into account robot constraints (e.g. limited FoV under eye-in-hand camera configuration), evaluates feasibility, and adds intermediate actions (e.g. look for objects outside view) to ensure executability through iterative search-and-refinement.
@@
VLM Planner. The VLM outputs a sequence of actions $\pi_{VLM}$, where each action consists of an action type (e.g. Pick) and a target object or location. As shown in Fig. 4(a), the prompt specifies the task goal, known locations, objects, and available actions, along with representation constraints and a fixed format to ensure a clear and ordered action plan. This sequence serves as a skeleton plan [56], which is subsequently grounded and refined. We utilize a foundation VLM [1] without fine-tuning to preserve generalization and maintain robust reasoning across diverse tasks.
@@
We divide $\pi_{VLM}$ into subtasks $\Pi=\{\pi^{1},\dots,\pi^{N}\}$ at placement actions, ensuring free robot hands for independent subtask execution. Each subtask labels object manipulation actions (e.g., Pick) as key actions $A_{k}$ and others (e.g., LookFor) as connecting actions $A_{c}$. Key actions are validated sequentially; if an action fails, a backward breadth-first search is triggered to find satisfying predecessor sequences using $A_{c}$ and the order-independent primitive actions $A_{s}$. This process includes actions to move the robot joints into suitable configurations when they are not appropriate for executing the next action. While performing search, it tracks the visited plans $V$ and records used actions $U$. Search failure returns unsatisfied preconditions and partial plans to VLM for replanning.
@@
Plan Refinement bridges VLM reasoning and robot execution through iterative feedback. Refinement occurs in two cases: (1) PDDL syntax errors during translation, generating format correction messages like ”Failed to create {action} instance: {error}.”; and (2) task planning errors during search when connecting actions cannot be found, prompting precondition error messages with partial plans. This process iterates until producing a fully grounded, executable plan, as shown Fig. 2.
@@ III-C Demonstration-Guided Robot Action Execution
The robot action execution stage implements the action sequence generated in Section III-B through the trajectory and mesh database from Section III-A. This execution framework integrates three core components: (1) target object pose estimation, (2) target trajectory generation, and (3) collision-aware motion planning.
Action Manager sequentially processes planner-generated actions by selecting target meshes and primitive skill, generating motions, and evaluating feasibility and termination. Target mesh selection maps VLM action parameters (e.g. cola_on_the_shelf) to corresponding mesh database objects (e.g. cola) using an LLM with simple prompts, as shown in Fig. 4(b). Action feasibility is evaluated based on predefined precondition that depend on the current robot and environment states. Upon action completion, the action manager notifies the task planner to dispatch the next action.
@@
During the second stage, the robot tracks $^{B}\tau_{\mathrm{al}}$ using the same QP-based IK solver with adaptive accuracy. The solver tolerance is gradually tightened as the end-effector approaches the target, prioritizing collision avoidance and joint limit handling during the early phase, and enforcing higher precision near task completion. This strategy enables safe obstacle avoidance during the approach while ensuring accurate manipulation near task completion [31].
@@ IV Experiments
The evaluation is conducted through real-world experiments using a Kinova Gen3 7-DOF arm equipped with a Robotis RH-P12-RN gripper and an eye-in-hand ZED stereo camera. For evaluation, we consider three sequential manipulation tasks (Fig. 6) using 9 different transparent and non-transparent objects (Fig. 7).
@@
Task2: Chemical Experiment. ”Can you make [target color] liquid in the [target object]?” This task simulates laboratory scenarios using transparent objects filled with various liquids. We tested with seven transparent objects and two target colors, requiring long-horizon planning, precise pouring, and collision avoidance in dense environments. Task3: Grocery Stocking. ”Arrange the [objects] on the staging area to the shelf in a straight row as shown in the reference image, placing each one directly in front of the previous one.” Inspired by retail scenarios, this automatic organization task requires long-horizon planning, collision avoidance, and precise pose estimation for aligned stocking.
For each task, we conducted 10 trials with different object instances and pose variations. The three single-object trajectories (pick, place, pour) were extracted from the object shown in Fig. 7-(1). Our framework supports 10 primitive actions (LookForAt, LookFor, Pick, Pour, PlaceBack, Place, PlaceBetween, PlaceInFront, Face, InitPose). The environments include a shelf (Task 1), a laboratory area (Task 2), and a staging area with a shelf (Task 3).
@@ IV-B Real World Results
Fig. 8 summarizes the task success rates and failure breakdowns for each task. We observed that all methods perform well on Tight Shelf Retrieval (Task 1), since most existing methods focus primarily on this type of task. However, for long-horizon tasks involving precise manipulation—such as chemical experiment and grocery stocking—the performance of YODO [52] and ClearGrasp [39] drops significantly.
@@
Figure 8: Comparison with state-of-the-art methods.
IV-C Ablation Studies
(1) Is depth estimation necessary for transparent object manipulation? To answer this question, we evaluate our method using raw depth sensor observations instead of our depth estimation results (Fig. 3). The comparison between Table I-(1) and Table I-Ours demonstrates that our depth estimation significantly improves performance across all tasks. Specifically, in the grocery stocking task, inaccurate depth information hindered the robot from compactly arranging objects within the shelf dimensions. This verifies the significant benefits of enhanced depth estimation for precise transparent robotics manipulation, as raw depth sensors alone are insufficient for handling transparent objects.
Table I: Ablation study of our core components on three different tasks in the real world.
@@ IV-C Ablation Studies
Depth 6D Pose Plan Refinement Plan Search Task 1 Task 2 Task 3
@@
(2) Is 6D pose estimation necessary compared to 3D position-based manipulation? One of the straightforward approaches for robot manipulation uses 3D position from open vocabulary segmentation and projects to 3D using depth for manipulation. As shown in Table I-(2), this position-based method achieves reasonable performance in Tight Shelf Retrieval (70%) and Chemical Experiment (60%) tasks, but fails in grocery stocking tasks, demonstrating the limitations of position-only approaches for grasping diverse rotation. Specifically, when objects are lying with small rotation errors, grasps often fail, highlighting the necessity of accurate 6D pose for robust manipulation.
(3-4) Can VLMs effectively handle long-horizon task planning? While naive VLMs may appear capable of long-horizon task planning, LLMs lack the ability to incorporate robot manipulation constraints into action sequences For example, in a pick-and-place task, the model may attempt to place an object before actually grasping it. Table I-(3) shows that a naive LLM mostly fails to generate valid plans, even for a short-horizon task such as tight shelf retrieval. An alternative approach is to iteratively refine the plan using a validator and repeated queries with error feedback, as in prior studies [60]. However, Table I-(4) demonstrates that this approach still struggles with long-horizon tasks, such as chemical experiments and grocery stocking, when the number of refinement iterations is limited to 10. Even when task planning succeeds, execution failures can still occur during the motion planning stage because this method relies solely on symbolic states without incorporating continuous variables such as joint poses, unlike our search-based method. Compared to Table I-(3-4) of VLM planning, our proposed search-based refinement successfully handles long-horizon planning.